OpenStudy (jazzyfa30):

5|2x - 5 | -3 = 12 please help

OpenStudy (anonymous):

5|2x - 5 | -3 = 12 --> 5|2x - 5 | = 15 --> |2x - 5 | = 3 --> Equation 1: 2x - 5 = -3 Equation 2:2x - 5 = 3

OpenStudy (anonymous):

Equation 1: 2x = 2 Equation 2:2x = 8

OpenStudy (anonymous):

Equation 1: x=1 Equation 2: x=4

OpenStudy (jazzyfa30):

it is neither one of those

OpenStudy (whpalmer4):

5|2(1)-5|-3 = 12 5|-3| -3 = 12 5*3 -3 = 12 5|2(4)-5|-3 = 12 5|8-5|-3 = 12 5(3)-3=12 those are both correct.

OpenStudy (whpalmer4):

but you are right in that it is neither one of those — it is both of those...

OpenStudy (jazzyfa30):

my problem says it is incorrect

OpenStudy (whpalmer4):

perhaps you didn't copy the problem correctly.

OpenStudy (tkhunny):

@jazzyfa30 I suggest you show some of YOUR work. That might settle the issue.

OpenStudy (jazzyfa30):

i copied and pasted it

OpenStudy (whpalmer4):

well, feel free to point out the mistake in my work, or FMP's :-)

OpenStudy (whpalmer4):

Mathematica agrees with our answers, too.

OpenStudy (jazzyfa30):

never mind ill figure it out

OpenStudy (whpalmer4):

Well, please post the correct answers for us when you do!